OpenStudy (anonymous):

(1/3x - 1/12y =2) (1/3x + 1/6y=-10)

OpenStudy (anonymous):

add them together, you get \(\frac{1}{12}y=8\)

OpenStudy (anonymous):

np

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

if the fractions annoy you, you are always free to multiply by any number you like for example you could multiply both sides of both equations by 12 and you would have no more fractions to deal with

OpenStudy (anonymous):

i will figure it just hte example in the book doesn't make sense thanks tho

OpenStudy (anonymous):

so i am taking 12 to multiply for both side ?

OpenStudy (anonymous):

yes that will clear the fractions for you

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

you get \[4x-y=24\]\[4x+2y=-120\]

OpenStudy (anonymous):

now you can subtract the second equation from the first, or else multiply the first equation by 2 and add

OpenStudy (anonymous):

ok thanks for your help sir or mam

OpenStudy (anonymous):

yw